OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0,6) and (1,3)

OpenStudy (anonymous):

Answer choices: (y - 3) = -3(x - 1) (y + 3) = 3(x + 1) (y + 3) = -3(x + 1) (y - 3) = 3(x + 6)

OpenStudy (anonymous):

@DSS @Sheraz12345 I need help on this one. Can you help me out on this one? :)

OpenStudy (sheraz12345):

a

OpenStudy (anonymous):

Thank you :)

OpenStudy (sheraz12345):

Np